OpenStudy (baldymcgee6):

if sin(x)=7/9 what is the value of tan(x)?

OpenStudy (baldymcgee6):

i got an answer of tan(x)=7/(4sqrt(2)) is this correct?

OpenStudy (unklerhaukus):

\[\sin x=\frac{\text{opposite}}{\text{hypotenuse}}\] \[\tan x=\frac{\text{opposite}}{\text{adjacent}}\]

OpenStudy (unklerhaukus):